OpenStudy (anonymous):

Let the sequence {an} be defined so that a1 = 1 and each succeeding term is found by multiplying the one before it by 4. Find the formula representing an, the nth term of the sequence.

OpenStudy (anonymous):

\[a _{n} = a _{n-1} \times4\]

jhonyy9 (jhonyy9):

a1=1 a2=4 a3=16 a4=64 . . . . an=an-1 x4 - yes this will be the nth term

OpenStudy (anonymous):

The non recursive solution would be:\[a_n = 4^{n-1}\]as a1 = 4^0 = 1, a2 = 4^1 = 4, a3 = 4^2 = 16, and so on.